# Quantum-enhanced sensing of spin-orbit coupling without fine-tuning

### v2

@@ 6 Optimal measurement
Figure 5: Comparison between QFI and CFI. (a) For a single-particle probe with $L{=}100$, the top curves correspond to ground state and bottom curves correspond to thermal state at $k_{B}T{=}0.01t$. (b) Many-body interacting probe with $L{=}6$ and $U{=}5t$.
@@
While the QFI gives the ultimate bound for precision, the optimal measurement basis is not unique. In our work, we find that the lattice version of the particle current operator [88, 89], defined as $\mathcal{I}=i\sum_{j,\sigma}(c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma}-c^{{\dagger}}_{j,\sigma}c_{j+1,\sigma})$, generates a suitable basis for estimation. The CFI measured in the eigenbasis of $\mathcal{I}$ matches with the corresponding QFI for single-particle probe, as shown in Fig. 5(a). Thermal energy exceeding the gap and interaction result in small discrepancy between the CFI and QFI for small $\alpha_{z}$ , as shown in Figs. 5 (a) and (b), respectively.
7 Conclusion
@@
Due to the long reaching impact of SOC on various fields ranging from condensed matter systems to emerging quantum technologies, it is of utmost importance to estimate the SOC strength with high precision. In this paper, we employ tools from quantum sensing to achieve quantum enhanced sensitivity for the estimation of Rashba SOC in one dimensional quantum wires. This enhancement is closely connected to gap closing feature of the system. Unlike conventional criticality based quantum sensors, in which the energy gap closing happens at the phase transition point, our probe achieves quantum advantage with Heisenberg precision across a wide range. Our analysis with both single-particle and many-body interacting probes establishes quantum advantage for all such parameters within characteristic temperature range. We have extended our results to multi-parameter sensing scheme where multiple SOC terms can be jointly estimated. In addition, we show that a simple particle current measurement can closely reach the ultimate precision limit.
@@
YB acknowledges support from National Natural Science Foundation of China (Grant No. 12404551) and the China Postdoctoral Science Foundation (Grant No. 2024M750339). AB acknowledges support from National Natural Science Foundation of China (Grants Nos. 12050410253, 92065115, and 12274059) and the Ministry of Science and Technology of China (Grant No. QNJ2021167001L). SS acknowledges support from National Natural Science Foundation of China (Grant No. W2433012).

### v3

@@ VI Optimal measurement
Figure 7: Comparison between QFI and CFI. The CFI is computed in the eigenbasis of the current operator defined in Eq. (16). (Top row) The dotted lines and the circles denote the QFI and CFI, respectively. (a) For a single-particle probe with L=100L{=}100, the top curves correspond to ground state and bottom curves correspond to thermal state at $k_{B}T{=}0.01t$. (b) Many-body interacting probe with L=6L{=}6 and U=5tU{=}5t. (Bottom row) The relative difference $\Delta\mathcal{F}$ between QFIM and CFIM across a range of $\alpha_{y}$ and $\alpha_{z}$ for L=100L{=}100 and B=0.01tB{=}0.01t. (c) The $(\alpha_{z},\alpha_{z})$ component of $\Delta\mathcal{F}$. (d) The $(\alpha_{y},\alpha_{y})$ component of $\Delta\mathcal{F}$. (e) The $(\alpha_{y},\alpha_{z})$ component of $\Delta\mathcal{F}$.
While the QFI gives the ultimate bound for precision, the optimal measurement basis is not unique. Although it is mathematically possible to find such a basis by diagonalizing the aforementioned SLD operator, this procedure usually leads to a complicated and non-local measurement choice. In practice, it is therefore advisable to look for an experimentally implementable measurement basis with which the calculated CFI is quite close to the QFI. In our work, we find that the lattice version of the particle current operator Hetényi (2009); Dutta et al. (2012), defined as
$\mathcal{I}=i\sum_{j,\sigma}(c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma}-c^{{\dagger}}_{j,\sigma}c_{j+1,\sigma}),$ (16)
@@
generates a suitable basis for estimation. The CFI measured in the eigenbasis of $\mathcal{I}$ matches with the corresponding QFI for single-particle probe, as shown in Fig. 7(a). In fact, by carrying out a Fourier transform to the momentum-space, we can readily see that $\mathcal{I}{=}\sum_{k}2\sin{(k)\,c^{{\dagger}}_{k,\sigma}c_{k,\sigma}}$. Therefore, the eigenbasis is the simple momentum basis which can be measured with existing experimental techniques such as the time-of-flight method. It is also possible to measure such current operators in optical lattice experiments Impertro et al. (2024, 2025). Thermal energy exceeding the gap and interaction results in small discrepancy between the CFI and QFI for small $\alpha_{z}$ , as shown in Figs. 7 (a) and (b), respectively.
The multi-parameter sensing scenario is more challenging because here the different optimal measurement bases for different parameters do not necessarily commute with each other Ragy et al. (2016) and one also needs to make sure that the matrices are not singular Mihailescu et al. (2025). Fortunately, for the model considered here, the current operator $\mathcal{I}$ provides the optimal basis for estimating both parameters as the CFIM $\mathcal{F}^{C}$ (defined in Eq. (2)) calculated in this basis matches quite well with the QFIM. We define the relative difference between the QFIM and CFIM by $(\Delta\mathcal{F})_{i,j}{=}(\mathcal{F}^{Q}{-}\mathcal{F}^{C})_{i,j}/\mathcal{F}^{Q}_{i,j}$. In Figs. 7 (c-e), we show that different elements of the $\Delta\mathcal{F}$ matrix are indeed very close to zero across a wide range of $\alpha_{z}$ and $\alpha_{y}$.
VII Precision analysis
Figure 8: Sensitivity from optimal SNR for system size L=100L{=}100. (a) Single-particle probe. (b) Many-body non-interacting probe with half-filling. Multi-parameter scenario with precision given by QFIM elements (c) $(\mathcal{F}^{Q})_{\alpha_{z},\alpha_{z}}$ and (d) $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{y}}$.
We now examine the precision obtained for the various probes considered in this work. The favorable scaling behavior of QFI without needing fine-tuning is quite encouraging, but we can also notice that the value of QFI itself can depend on the parameters themselves. Since the QFI exhibits a nearly consistent super-linear scaling $\sim L^{\beta}$ across a broad parameter range, its dependence on the unknown parameter comes from the prefactor. This can be seen, for example, in Figs. 1(d), 2(b), and in Fig. 6. It is quite evident from Fig. 1(c), that for a given $B$ value, the maximum QFI occurs at a certain $\alpha_{z}$. Therefore, it is crucial to analyze the sensitivity across the ranges of the parameters to examine the merit of the metrological advantage. To this effect, we report the SNR in Fig. 8 which shows the ratio of the parameter and the precision achievable in the optimal case. As mentioned in Sec. II.3, this is given by $\alpha_{z}\sqrt{MF^{Q}(\alpha_{z})}$ for the single-parameter case, and by $\alpha_{i}\sqrt{M/(\mathcal{F}^{Q})^{-1}_{i,i}}$ for the multi-parameter case. We observe that, even with experimentally low-cost value measurement repetition number M=1000M{=}1000, the SNR values obtained are quite high, as SNR in the range 1010010-100 is already considered satisfactory in physical realizations. This is shown for single-particle probes in Fig. 8(a), for many-body non-interacting probes in Fig. 8(b), and for the multi-parameter scenario in Figs. 8(c) and (d). As is evident from the Figs. 8(a)-(b), the SNR values are lower for small $\alpha_{z}$ values, but the estimation errors are still at least one order of magnitude smaller than the parameter values themselves. This relays the importance of the SNR analysis. We also notice that the SNR values create certain arrays of local maxima for the many-body case (Fig. 8(b)), for specific combinations of $\alpha_{z}$ and $B$ values. This is caused by the overall complicated spectrum of the filled energy levels that results in local minimum in energy gap structure, which subsequently increases the QFI at those points.
Beyond optimal estimation, one can carry out similar analysis by using more practical estimation strategies such as Bayesian approaches. We show these results in Fig. 9 by quantifying the sensitivity as the ratio of the estimated parameter and the achieved precision for the same number of measurement repetitions as before. In this case, we utilize the basis laid out in Sec. VI to generate the measurement data. The measurement data is generated by sampling from the probability distribution provided by the probe state in this basis. The posterior distribution obtained, as outlined in Sec. II.2, provides the error that quantifies the sensitivity. Again we observe quite satisfactory values for the sensitivity obtained for the single particle probe in Fig. 9(a) and a many-body interacting probe in Fig. 9(b). Therefore we can infer that the performances of the different probes are quite competent, along with the advantages coming from the scaling properties and the lack of the requirement of fine-tuning.
Figure 9: Sensitivity from Bayesian analysis. (a) Single-particle probe with L=100L{=}100. (b) Many-body interacting probe with L=6L{=}6 and U=5tU{=}5t.
VIII Conclusion
@@
Due to the long reaching impact of SOC on various fields ranging from condensed matter systems to emerging quantum technologies, it is of utmost importance to estimate the SOC strength with high precision. In this paper, we employ tools from quantum sensing to achieve quantum enhanced sensitivity for the estimation of Rashba SOC in one dimensional quantum wires. This enhancement is closely connected to gap closing feature of the system. Unlike conventional criticality based quantum sensors, in which the energy gap closing happens at the phase transition point, our probe achieves quantum advantage with Heisenberg precision across a wide range. Our analysis with both single-particle and many-body interacting probes establishes quantum advantage for all such parameters within characteristic temperature range. We have extended our results to multi-parameter sensing scheme where multiple SOC terms can be jointly estimated. In addition, we show that a simple particle current measurement can closely reach the ultimate precision limit. This work can be extended to search for other models where enhanced sensitivity can be achieved over a wide range. Parallelly, one can also extend the study for other types of SOC and even other types sensors such as field sensors. Another avenue worth exploring is the possibility to utilize non-equilibrium dynamics for estimating SOC as they are experimentally more stratightforward to employ.
Acknowledgements.
@@
SS acknowledges support from National Natural Science Foundation of China (Grant No. W2433012). AB acknowledges support from the National Natural Science Foundation of China (grants No. 12274059, No. 12574528, No. 1251101297 and No. W2541020).
Data availability
The data that support the findings of this article are openly available git .